OpenStudy (anonymous):

Solve the following initial value problems: y'+2y = 2+e^x; y(0)=4 thnx in advance :)

terenzreignz (terenzreignz):

Let's summon the Mimi.... @Mimi_x3

OpenStudy (anonymous):

mimi i choose you :P

terenzreignz (terenzreignz):

okay, in lieu of Mimi, you'll have to work with me :P

OpenStudy (anonymous):

ok :D sounds good

terenzreignz (terenzreignz):

Any idea what method we're supposed to be using?

OpenStudy (anonymous):

i think we have to find C

OpenStudy (anonymous):

then subtitute into the simplified equation, but i still cant quite get it xD

terenzreignz (terenzreignz):

First we need to solve this differential equation :D

OpenStudy (anonymous):

ok lets start :)

terenzreignz (terenzreignz):

This needs what's called an integrating factor.

terenzreignz (terenzreignz):

Pikachu's here.... @Mimi_x3 ?

OpenStudy (anonymous):

xD

terenzreignz (terenzreignz):

Okay, let's write that equation down... \[\Large y' +2y = 2+e^x\]

terenzreignz (terenzreignz):

Whenever you have a differential equation of the form... \[\Large y' + p(\color{red}x)\color{blue}y=q(\color{red}x) \] You need to find what's called an integrating factor.

terenzreignz (terenzreignz):

What you do is take this part... \[\Large y' + \boxed{p(\color{red}x)}\color{blue}y=q(\color{red}x)\] and integrate it.

OpenStudy (anonymous):

the 2?

terenzreignz (terenzreignz):

Very good! \[\Large y' +\boxed2y = 2+e^x\] so I want you to find \[\Large \int 2 \ dx\] Shouldn't be too hard, really :D

OpenStudy (anonymous):

so it would be 2y^2/2?

terenzreignz (terenzreignz):

huh? No, first the integral, can you solve it? :D \[\Large \int 2 \ dx\]

OpenStudy (anonymous):

2x

terenzreignz (terenzreignz):

Okay. So the integrating factor is \[\Large e^{2x}\]

terenzreignz (terenzreignz):

We multiply this to the entire expression... \[\Large e^{2x}\frac{dy}{dx}+2e^{2x}y=2e^{2x}+e^{3x}\] Catch me so far?

OpenStudy (anonymous):

yes, but why did we do that? sorry for the dumb question

terenzreignz (terenzreignz):

Very good question. (not a dumb one ^.^)

terenzreignz (terenzreignz):

You know implicit differentiation?

OpenStudy (anonymous):

hmmm took it last sem but i kinda forgot, need a reminder :/

terenzreignz (terenzreignz):

Okay, you notice, if we differentiate this implicitly \[\Large y e^{2x}\] Using the product and chain rules, we get \[\Large e^{2x}y'+2e^{2x}y\] Right?

OpenStudy (anonymous):

hmm

OpenStudy (anonymous):

oh yea ok got it

terenzreignz (terenzreignz):

Well, that's what our left-side is now. \[\Large \color{red}{e^{2x}\frac{dy}{dx}+2e^{2x}y}=2e^{2x}+e^{3x}\]

OpenStudy (anonymous):

y' = dy/dx yep

terenzreignz (terenzreignz):

Okay, so the entire left side may NOW be written \[\Large \frac{d}{dx}(ye^{2x})=2e^{2x}+e^{3x}\]

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

you may continue :3 if you want

terenzreignz (terenzreignz):

Sorry. It's now basically a separable differential equation. \[\Large d(ye^{2x}) =( 2e^{2x}+e^{3x})dx\]

OpenStudy (anonymous):

yea we moved the dx to the RHS :) nice

OpenStudy (anonymous):

we intigrate now?

terenzreignz (terenzreignz):

Yup. Left side just integrates simply. \[\Large ye^{2x} =\int( 2e^{2x}+e^{3x})dx\] right side... you do that.

OpenStudy (anonymous):

we integrate 2e^2x then e^3x... (2e^2x+1/2) + (e^3x+1/2) + C ? gah i think its wrong :/

terenzreignz (terenzreignz):

Just relax... deep breaths and try again :)

OpenStudy (anonymous):

2e^2x+1 is that right? O.o

terenzreignz (terenzreignz):

ahaha... this is an exponential, not a polynomial :D

OpenStudy (anonymous):

hmmmm, where did i go wrong on my prev ans xD

terenzreignz (terenzreignz):

Remember that \[\Large \int e^xdx = e^x+C\]

OpenStudy (anonymous):

ooh yea

OpenStudy (anonymous):

sry i forgot that one

terenzreignz (terenzreignz):

So redo it :)

OpenStudy (anonymous):

so its the same we just add +c to the end?

terenzreignz (terenzreignz):

Not simply. It requires a little substitution.

OpenStudy (anonymous):

am out of ideas...

terenzreignz (terenzreignz):

You really should practice Integrals... THEY ARE ESSENTIAL to differential equations :D

terenzreignz (terenzreignz):

I'll integrate for you this once :D \[\Large ye^{2x} =e^{2x}+\frac13e^{3x}+C\]

terenzreignz (terenzreignz):

So we need to find C.

OpenStudy (anonymous):

we use x = 0 and y = 4 right?

terenzreignz (terenzreignz):

Yup.

OpenStudy (anonymous):

do we need to simplify further or just substitute?

terenzreignz (terenzreignz):

Just substitute. \[\Large 4(e^{2(0)}) = e^{2(0)}+\frac13e^{3(0)}+C\] And solve for C.

OpenStudy (anonymous):

4= 1.3+c

terenzreignz (terenzreignz):

No, don't use decimals, use fractions.

OpenStudy (anonymous):

2.7 = c?

OpenStudy (anonymous):

lol sry

terenzreignz (terenzreignz):

and it's 4 = 1 + (1/3) + C

terenzreignz (terenzreignz):

(1/3) is not 0.3

OpenStudy (anonymous):

o_o ok

OpenStudy (anonymous):

i think i should change my major

terenzreignz (terenzreignz):

So, just solve this \[\Large 4 = 1+\frac13 + C\] I have to go for now.

terenzreignz (terenzreignz):

I'll be back in 20 minutes, maybe :D

OpenStudy (anonymous):

ok take your time

OpenStudy (anonymous):

aha ok thank you

OpenStudy (anonymous):

c = 8/3

OpenStudy (anonymous):

its ok, thanks for the suggestion

OpenStudy (anonymous):

haha why are you worrying too much? isn't this site for multiple opinions?

terenzreignz (terenzreignz):

There's virtually no such thing as opinions in Maths. Either I'm correct or I'm wrong.

OpenStudy (anonymous):

there are ways to get to the answer right?

terenzreignz (terenzreignz):

Surely.

OpenStudy (anonymous):

oh and btw sry i meant c= 16/3 just a little miscalculation

OpenStudy (anonymous):

then we subtitute into the equation's C with 16/3 right?